The outstanding capital stock of Flounder Corporation consists of 2,000 shares of $100 par value, 8% preferred, and 4,500 shares of $50 par value common.

Assuming that the company has retained earnings of $90,500, all of which is to be paid out in dividends, and that preferred dividends were not paid during the 2 years preceding the current year, state how much each class of stock should receive under each of the following conditions.

(a) The preferred stock is noncumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places


(b) The preferred stock is cumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places


(c) The preferred stock is cumulative and participating. (Round the rate of participation to 4 decimal places, e.g.1.4278%. Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places
Answer #1

PageNo Answer: Given that, => pore ferved shore dividends = 2000 x 100 x 87. - 2000 X 100 X0.8 = 16,000 - Common share dividePageNo@ b) The preferred stock iş Cumulative and nonparticipating Preterred = 16,000 X3 = 48,000 Common = (90.500 -48,000) 42Pagento Common stock = 4,500X 50 = 225,000 Ratio of participation Poreferred - 200,000 Ua5,000 = 47.05884. Common stock - 225Page No Poreferred = 48,000 + 11,529 = 59,529 Common = 18,000 + 12,941 = 30,941